OpenStudy (anonymous):

Tori works two jobs to pay for college. She tutors for $30 per hour and also works as a receptionist for $10 per hour. Due to her class and study schedule, Tori is only able to work up to 20 hours per week, but must earn at least $200 per week. If t represents the number of hours Tori tutors and r represents the number of hours she works as a receptionist, which system of inequalities represents this scenario

OpenStudy (anonymous):

well : t= $30/ h ( as tutor) r= $10/h (as receptionist) let h (t) = hours worked as tutor h (r) = hours worked as receptionist 1) now she can work MAX 20 h / week => h (t) + h (r) < = ( means less or equal to ) 20 ( note that u add : hour + hour = hour) 2) must earn at least $200/ week => t * h (t) + r * h (r) >= ( more or equal to ) $200 or $30/ h * h (t) + $10/h * h ( r) >= ( more or equal to ) $200

OpenStudy (anonymous):

if math has such problims why r u so good with it